# Forests of Uncertaint(r)ees: Using tree-based ensembles to estimate probability distributions of future conflict

### v1

@@ 4.1 Aggregate Performance
To evaluate the overall performance of our three ensemble models [^12] For simplicity, we subsequently use the term “model” in this context to refer to the respective specification used to generate our global-only, local-only, or global-local ensemble prediction., we compare them to several benchmarks provided by the VIEWS team across the six yearly test windows. Those are two naive benchmarks and three “conflictology” benchmarks based on medium- to long-term conflict history. The naive benchmarks are samples drawn from a Poisson distribution centered around the last observed values for each grid cell and predictions with only zero values. The “conflictology” benchmarks all treat historic fatality counts as draws from the predictive distribution to generate forecasts. The first benchmark (“conflictology”) uses fatality counts from a specific grid cell during the previous 12 months for the respective prediction window (12 draws). The second benchmark (“conflictology neighbors”) follows the same principle, but uses the combined conflict history of the grid cell and its immediate neighbors (108 draws). The third benchmark (“bootstrap 240”) draws 1000 random samples from the grid cell’s conflict history of the last 240 months. All also adhere to the two-month gap between training and test data. For instance, the “conflictology” samples for all months in 2018 are based on the observed fatalities from November 2016 to October 2017 (Hegre et al., 2025). [^13] Information in addition to Hegre et al. (2025) based on the source code at https://github.com/prio-data/prediction_competition_2023/blob/a45796ce8d1ffdd82e879e05c46d90c58b460a66/benchmark.py.
The evaluation results for our models and the benchmarks are reported in Table 1. We base our evaluation on the challenge’s main metric CRPS and the two additional metrics specified in the prediction challenge, the Ignorance Score (IGN) and the Mean Interval Score (MIS), all of which evaluate samples from a distribution (Hegre et al., 2025). For all three scores, lower values are better and perfect predictions have a score of zero. For potential comparisons with point prediction models, we also report Mean Squared Error (MSE) and Mean Absolute Error (MAE) values based on point predictions calculated via maximum a posteriori (MAP) estimation. We use Gaussian kernel density estimation (KDE) to estimate the probability density function (PDF) for each sample, through which we select the maximum.
At first glance, our three models are reliably able to beat the benchmarks, with very few exceptions. Our global model performs slightly better than the local model, while the global-local combination performs about the same as the global model. Depending on the metric, it even pulls ahead in two to three out of the six test windows. While this means we are unable to exploit systematic differences across contexts with our approach to improve predictive performance, this still opens avenues for the selective inclusion of data with limited geographic availability.
@@
Table 1: Overview of model and benchmark metrics. Best results for each year are marked in bold. Lower scores signify better performance. Note that the CRPS is always equal to the MAE in the case of all-zero predictions.
@@ 4.2 Uncertainty in Model Evaluation
At second glance, differences in performance between both our own models and our models’ improvements over most benchmarks are miniscule, and could conceivably be caused e.g. by random components in the modeling process, warranting a closer look at the scoring functions and their properties. First, it should be noted that all three scores cannot be compared across different data and thus years. Yearly scores are calculated as the mean of the score for all $N$ pgms (Hegre et al., 2025) [^14] See https://github.com/prio-data/prediction_competition_2023 for the challenge-specific implementations of the MIS and IGN scores., e.g.
$\overline{CRPS}=\frac{1}{N}\sum_{i=1}^{N}CRPS(F_{i},y_{i})$ (1)
with
$CRPS(F_{i},y_{i})=\int_{-\infty}^{\infty}\left(F_{i}(x)-\mathbbm{1}(x-y_{i})\right)^{2}\,dx\quad\text{where}\quad\mathbbm{1}(z)=\begin{cases}1&\text{if }z\geq 0\\ 0&\text{otherwise}\end{cases}$ (2)
for the PDF $F_{i}(x)$ representing the prediction distribution and the corresponding observation $y_{i}$.
Both $CRPS_{i}$ and $MIS_{i}$ for all-zero predictions and no observed violence are 0, while $IGN_{i}$ is a fixed constant. Assuming most zeroes are predicted correctly, the highly zero-inflated nature of our prediction task means the informative range for each score becomes very small, as only a tiny percentage of predictions actually impact the sum, while N remains constant. This means the scores are not independent of the underlying data for a given application. For illustration, CRPS and MIS scores are almost perfectly correlated with the number fatalities across the different years, while IGN scores are strongly correlated with the number of non-zero pgms for most models and benchmarks (Table 2).
@@
Table 2: Correlation between evaluation scores and number of fatalities / number of non-zero pgms across the six test windows for global model and benchmarks. Note that the perfect correlation for all-zero samples with fatalities for CRPS and MIS, and with non-zero pgms for IGN is expected behavior. The perfect correlation of the Bootstrap 240 benchmarks with fatalities for the CRPS is the result of rounding, while for the MIS it is the result of the 90% prediction interval used (Hegre et al., 2025) in combination with the composition of the individual samples. As the samples never contain more than 1.9% non-zero draws for this benchmark in a given test window, the intervals only contain all-zero predictions.
@@
Figure 3: Value ranges of CRPS scores for one year of simulated actuals and predictions, with varying noise and “accuracy”. Actuals are simulated with zero-inflatedness matching the training data, and values drawn from an estimated PDF based on non-zero observations < 1000 fatalities. Predictions are drawn from a poisson distribution with mean and variance equal to the corresponding actual. “Accuracy” ($\alpha$) less than 1 means a share of non-zero samples was replaced with all-zero predictions. Noise ($n$) means actuals were shifted by $50n\varepsilon$ [^15] While the cutoff point for the noise multiplier is chosen arbitrarily, the range should be sufficient to produce fairly strong effects given the values seen in our predictions - based on Figure 1 and a median number of fatalities of 5 in the training data. Increasing the multiplier much beyond $50\varepsilon$ continues to decrease the size of the CRPS change when varying the “accuracy”.before creating the samples, with $\varepsilon$ randomly drawn from a uniform distribution between -1 and 1.
@@
Given this, it is essential to understand the informative range of the score to properly evaluate our models. We use simulated data to experimentally explore the impact of prediction accuracy for this degree of zero-inflation. For this, we focus on the CRPS as the main metric of the challenge. For our simulated observations $z_{i}$, we create a sample matching the size of one test window (N=157.320) and the share of non-zero pgms from the dataset (0.005, N=787). We draw simulated non-zero values $z_{i}^{+}$ based on a PDF estimated via Gaussian KDE from all observed values $x_{i}$ in the training data, where $0<x_{i}<1000$, thus controlling for extreme outliers, multiplying all resulting negative values with -1 to ensure we have valid fatality numbers. Next, simulated predictions are created based on these simulated observations. “Perfect” predictive distributions are created by drawing samples with size N=1000 from a Poisson distribution, setting the expectation $\lambda$ to the simulated actual value, i.e. $\lambda_{i}=z_{i}$ (see Hegre et al., 2023), thus obtaining a predictive distribution for each simulated observation $z_{i}$. Finally, we introduce errors to our predictions in two ways: First, we vary the “accuracy” $\alpha$ by replacing a share of the predictions for the simulated non-zero observations $z_{i}^{+}$ with all-zero predictions, i.e. for $\alpha=0.8$, 20% (157) of the 787 predictions for simulated non-zero observations are replaced with all-zero predictions. Second, we add different levels of random noise to all simulated non-zero observations before drawing the samples for our predictive distributions, i.e. $z_{i}^{+\prime}=z_{i}^{+}+\varepsilon$, randomly shifting the center of our predictive distributions away from the actual value. In both cases we still evaluate against the original $z_{i}$. [^16] We replace any values ¡ 0 with 0 before drawing the samples, to ensure the prediction samples only contain non-negative values.
@@
Figure 15 shows the CRPS value ranges based on the simulation across varying “accuracies” and noise. CRPS values range from 0.004 to .101 going from “perfect” predictions to 0.1 “accuracy”, dropping on average .01 per 10% drop in “accuracy”. The more noise is added to the data, the smaller the change in CRPS becomes when varying the “accuracy”, with the changes more than half for the maximum level of noise in our simulation. Given that our best models are on average 0.007 points better than the best benchmark, this makes us confident that this means a real improvement in performance over the benchmarks rather than random variation.
@@
Figure 4: Comparison of model-wide scores with scores based on grid cells grouped by individual countries for the three distributional metrics: CRPS (a), MIS (b), and IGN (c).
@@
Finally, having confirmed that the small differences in our evaluation metrics matter, we want to understand where these differences come from. To do so, we match grid cells to countries based on a majority rule and evaluate our global model for each country individually. Country borders for assignment are taken from geoBoundaries (Runfola et al., 2020). Since this approach means strongly varying levels of zero-inflatedness across the countries, the corresponding scores also vary strongly and can no longer be directly compared (Figure 4). Instead, we rank the performance of all models and benchmarks for each country and calculate the average ranks across all scenarios (Table 3). Based on mean rank across all countries and years, our model comes in only 5th on the CRPS, 4th on the MIS, and 3rd on the IGN. This can be explained by roughly half of all countries never experiencing any violence, for which our model gets heavily penalized in a rank-based evaluation due to its base assumption of a small but non-zero chance of violence anywhere. [^17] For both the CRPS and the IGN, any non-zero values in our samples mean that predictions will be scored marginally worse. Since the benchmarks generally yield all-zero samples in countries without prior violence, this means our model will be ranked 6th in such cases, while the benchmarks all receive a joint 1st place. To illustrate: Out of 80 countries covered by our model, 32 never see any fatalities across all yearly test windows, with a minimum of 39 no-violence countries in any individual window. However, if we focus only on countries for which any violence at all was observed in a given year - what we argue is where the performance of our model matters most, our model ranks highest on average compared to all benchmarks and for all three distributional metrics.
@@ 5 Discussion
We demonstrate that it is possible to produce high-resolution conflict predictions with uncertainty estimates using a modeling architecture built from relatively simple components, which are able to beat a range of heuristic conflict benchmarks. We further find that these performance gains arise from cases where violence occurs, and thus in situations in which prediction tools are also most likely valuable to practitioners. With our evaluation, we highlight the need to not simply pursue marginal improvements in evaluation metrics, but to understand how they behave in the context of the prediction problem. While this is a useful first step towards actionable uncertainty estimates, the prediction intervals we produce are still fairly wide, with further research needed both to sharpen predictions and to reduce the sources of uncertainty.
@@
Building on our approach, data-related avenues for improvement naturally include the use of additional data sources, potentially in combination with multi-level modeling to combine data at different resolutions, while also accounting for the distinction between macro- and micro processes of conflict (Balcells and Justino, 2014; Fritz et al., 2024). As our results show, it is also possible to combine multiple regional models without significantly losing performance. This opens up avenues to integrate data sources previously discarded due to coverage issues, e.g. collected by regional organizations such as ECOWAS. Additionally, data cleaning techniques such as outlier correction, applied both to the target and to the features, could reduce data-driven uncertainties. In addition to general improvements through better-suited methods, approaches such as selective classification (El-Yaniv and Wiener, 2010; Hüllermeier and Waegeman, 2021) and out-of-distribution detection (Yang et al., 2024; Gruber et al., 2025) could be used on the modeling side to reject predictions where the estimated uncertainty is too high for the predictions to be useful in practice.
@@
Finally, as little is known about the composition of uncertainty in conflict modeling, more research is also needed to understand to what extent this is due to data issues and to what extent violence is the consequence of inherently stochastic processes. Much of this question relates back to conflict datasets and the challenges of recording events globally as discussed above. For example, fine-grained local datasets may report many instances of violence not contained in the larger data collection efforts needed for forecasting systems (Bazzi et al., 2022). In combination with differences in inclusion criteria and coding procedures (Raleigh et al., 2023; Öberg and Yilmaz, 2025), this leads to a large number of non-overlapping events among conflict event datasets based on very similar sources (Donnay et al., 2019), and makes it hard to establish a ground truth for further examination. Addressing this issue would go a long way towards understanding and quantifying the individual components of uncertainty in conflict forecasting, and thus enable a more targeted approach to its reduction.

### v2

@@ 4.1 Aggregate Performance
Table 1: Average model and benchmark metrics across all six yearly test windows. Best results for each year are marked in bold. Lower scores signify better performance. Note that the CRPS is always equal to the MAE in the case of all-zero predictions.
To evaluate the overall performance of our three models, we compare them to several benchmarks provided by the VIEWS team across the six yearly test windows. Those are two naive benchmarks and three “conflictology” benchmarks based on medium- to long-term conflict history. The naive benchmarks are samples drawn from a Poisson distribution centered around the last observed values for each grid cell and predictions with only zero values. The “conflictology” benchmarks all treat historic fatality counts as draws from the predictive distribution to generate forecasts. The first benchmark (“conflictology”) uses fatality counts from a specific grid cell during the previous 12 months for the respective prediction window (12 draws). The second benchmark (“conflictology neighbors”) follows the same principle, but uses the combined conflict history of the grid cell and its immediate neighbors (108 draws). The third benchmark (“bootstrap 240”) combines $1,000$ random draws from the grid cell’s conflict history of the last 240 months. All also adhere to the two-month gap between training and test data. For instance, the “conflictology” samples for all months in 2018 are based on the observed fatalities from November 2016 to October 2017 (Hegre et al., 2025). [^13] Information in addition to Hegre et al. (2025) based on the source code at https://github.com/prio-data/prediction_competition_2023/blob/a45796ce8d1ffdd82e879e05c46d90c58b460a66/benchmark.py.
We base our evaluation on the challenge’s main metric, the CRPS, and the two additional metrics specified in the prediction challenge: an adjusted, binned version of the Ignorance Score (IGN) (also called Log Score) and the Mean Interval Score (MIS). [^14] For consistency with Hegre et al. (2025), we follow the convention of referring to the Mean Interval Score rather than simply the Interval Score in the text, denoting that it is averaged across multiple pgms for the evaluation, while other scores are referred to without this prefix despite using the same averaging procedure (see equation 4 All scores evaluate samples from a distribution, with lower scores representing better predictions and perfect predictions corresponding to a score of zero for CRPS and MIS, and 0.014 for IGN (see equations 5 - 7).
For an observation $y_{i}$ and its corresponding predictive distribution $F_{i}(.)$, the CRPS is defined as
$CRPS(F_{i},y_{i})=\int_{-\infty}^{\infty}\left(F_{i}(x)-\mathbbm{1}(x-y_{i})\right)^{2}\,dx,\quad\text{where}\quad\mathbbm{1}(z)=\begin{cases}1&\text{if }z\geq 0\\ 0&\text{otherwise}\end{cases}\text{.}$ (1)
Following VIEWS, the adjusted, binned IGN is defined as
$ab\text{-}IGN(F_{i},y_{i})=-\log_{2}\left(\frac{n_{b_{i}}+\omega}{N_{sample}+N_{bins}\,\omega}\right),$ (2)
where $n_{b_{i}}$ denotes the count of draws in bin $b_{i}\in\mathcal{B}$ with $y_{i}\in b_{i}$, $N_{sample}$ is the total number of draws from $F_{i}$$N_{bins}$ is the number of individual bins, and $\omega$ is a constant adjustment value. In line with the challenge, we use $N_{sample}=1,000$, $\omega=1$, and the bins $\mathcal{B}=\{\{0\},[1,2],[3,5],[6,10],[11,25],[26,50],[51,100],$ $101,250],[251,500],[501,1000],[1001,\infty)\}$, i.e. $N_{bins}=11$.
The Interval Score (IS) is defined as
$IS(F_{i},y_{i})=(U_{i}-L_{i})+\frac{2}{q}(L_{i}-y_{i})\times\mathbbm{1}(L_{i}>y_{i})+\frac{2}{q}(y_{i}-U_{i})\times\mathbbm{1}(y_{i}>U_{i}),$ (3)
with $U_{i}$ denoting the upper $1-\frac{q}{2}$ quantile, $L_{i}$ denoting the lower $\frac{q}{2}$ quantile of $F_{i}$, and the same indicator function $\mathbbm{1}$ as in (1). The challenge uses a 90% prediction interval, which means q=0.1q=0.1.151515CRPS is calculated using the xskillscore Python package (Bell et al., 2021). See https://github.com/prio-data/prediction_competition_2023 for the challenge-specific implementations of the MIS and IGN scores.
Yearly scores for all metrics are calculated as the mean of the score over all incuded pgms (Hegre et al., 2025), for example
$\overline{CRPS}=\frac{1}{N}\sum_{i=1}^{N}CRPS(F_{i},y_{i}).$ (4)
For potential comparisons with point prediction models, we also report Mean Squared Error (MSE) and Mean Absolute Error (MAE) values based on MAP estimates. We use Gaussian kernel density estimation (KDE) to estimate the probability density function (PDF), from which we select the maximum.
The evaluation results for our models and the benchmarks are reported in Table 1. Across all test windows, our three models are reliably able to beat the benchmarks, with very few exceptions. Our global and global-local models perform about the same, with the latter even pulling ahead in two to three out of the six test windows, depending on the metric. The local model is only slightly worse. While this means we are unable to exploit systematic differences across conflict contexts to improve predictive performance, this still opens avenues for the selective inclusion of data with limited geographic availability. The strong performance of our models is confirmed by the VIEWS team’s evaluation of the true future forecasts, in which our models ranked first among all contributions at the pgm level (VIEWS, 2026).
@@ 4.2 Uncertainty in Model Evaluation
Table 2: Correlation between evaluation scores and number of fatalities / number of non-zero pgms across the six test windows for the global model and the benchmarks. Note that the perfect correlation for all-zero samples with fatalities for CRPS and MIS, and with non-zero pgms for IGN is expected behavior. The perfect correlation of the Bootstrap 240 benchmarks with fatalities for the CRPS is the result of rounding, while for the MIS it is the result of the 90% prediction interval used (Hegre et al., 2025) in combination with the composition of the individual samples. As the samples never contain more than 1.9% non-zero draws for this benchmark in a given test window, the intervals only contain all-zero predictions.
Taking a closer look, differences in performance between both our own models and our models’ improvements over most benchmarks are miniscule, and could conceivably be caused by random components in the modeling process or simply be irrelevant in practice. This warrants a deeper examination of the scoring functions and their properties. First, it should be noted that all three scores in aggregate cannot be compared across the different years, since they are not dependent on the distributions of the underlying data. Furthermore, for the special case of an all-zero forecast, i.e. a forecast drawn from a predictive distribution $F_{i}^{0}$ that places all probability mass on zero $P(Y_{i}=0)=1$, with $F_{i}^{0}(0)=1$, and the highly-common $y_{i}=0$ observation, individual scores (1), (2), and (3) result in
$CRPS(F_{i}^{0},0)=0$ (5)
for the CRPS,
$ab\text{-}IGN(F_{i}^{0},0)=-\log_{2}(\frac{1001}{1011})\approx 0.014,\text{ since }n_{b_{i}}=1000\text{ with }y_{i}=0$ (6)
for the adjusted, binned IGN, and
$IS(F_{i}^{0},0)=0$ (7)
for the MIS, respectively.
The highly zero-inflated nature of our prediction task means only a tiny percentage of predictions actually impact the aggregate scores when the means across all relevant pgms are calculated (4), as $N$ remains constant. Assuming most zeroes are predicted "correctly" with all-zero forecasts, the low individual scores (5 - 7) correspondingly result in very low mean scores, with the informative range for each score becoming very small. Illustrating this effect, mean CRPS and MIS scores are almost perfectly correlated with the number fatalities across the different years, while mean IGN scores are strongly correlated with the number of non-zero pgms for most models and benchmarks (Table 2).
@@
Figure 3: Value ranges of CRPS scores with varying noise and “accuracy” based on simulated data. Simulated observations are created with zero-inflatedness matching the six test windows and values drawn from an estimated PDF based on real non-zero observations <1,000<1,000 fatalities. Simulated predictions are drawn from a Poisson distribution with mean and variance equal to the corresponding simulated observation. “Accuracy” ($\alpha$) <1<1 means a share $1-\alpha$ of non-zero predictions was switched with all-zero predictions. The Noise scale factor ($\sigma$) means observations were shifted by $\varepsilon=50\sigma u$ before creating the samples, with 1<u<1-1<u<1 drawn from a uniform distribution. The dark planes represents the scores for our best model in the respective test window.
@@
Given this, it is essential to understand the informative ranges of the scores to properly evaluate our models. We use simulated data to experimentally explore the impact of prediction accuracy for the degree of zero-inflation across all test windows. For our simulated observations $z_{i}$, we create samples matching the size (N=157,320N=157{,}320) and the share $p$ of non-zero pgms ($0.0078<p<0.0107$, corresponding to $1,277<N^{+}<1{,}683$ pgms) for each year. We draw the simulated non-zero values $z_{i}^{+}$ based on a PDF estimated via Gaussian KDE from all observed values $0<y_{i}<1,000$ in the training data, thus controlling for extreme outliers. We multiply all negative values in this sample with -1 to ensure valid fatality numbers. Next, simulated predictions are created based on these simulated observations. “Perfect” predictive distributions are created for each simulated observation $z_{i}$ by drawing samples with size $N_{sample}=1,000$ from a Poisson distribution, setting the expectation $\lambda$ to the simulated actual value, i.e. $\lambda_{i}=z_{i}$ (see Hegre et al., 2023). Finally, we introduce errors to these predictions in two ways: First, we vary the “accuracy” $\alpha$ by randomly swapping a share ($1-\alpha$) of the predictions for the simulated non-zero observations $z_{i}^{+}$ with other all-zero predictions, thus keeping the model calibration consistent. For example, for $\alpha=0.8$ in the 2018 test window, 20% (255) of the 1,277 prediction samples for simulated non-zero observations are replaced with all-zero predictions, and in turn inserted in the place of random all-zero samples. Second, we vary $\sigma$ to add different levels of random noise $\varepsilon=50\sigma u\text{, where }u\sim\mathcal{U}(-1,1)\text{ and }0\leq\sigma\leq 1$ to all simulated non-zero observations before drawing the prediction samples, [^16] While the cutoff point for the noise multiplier is chosen arbitrarily, the range should be sufficient to produce fairly strong effects given the values seen in our predictions – based on Figure 1 and a median number of fatalities of 5 in the training data. Increasing the multiplier much beyond $50\varepsilon$ continues to decrease the size of the metrics’ change when varying the “accuracy”. i.e. $z_{i}^{+\prime}=z_{i}^{+}+\varepsilon$, randomly shifting the center of our predictive distributions away from the actual value. [^17] We replace any values ¡ 0 with 0 before drawing the samples, to ensure the prediction samples only contain non-negative values. In both cases we still evaluate against the original $z_{i}$.
@@
Figure 3 shows the simulated CRPS value ranges for the share of non-zero observations across all test windows and across varying accuracies and noise levels in relation to our best model. Except for 2021 and 2022, where outliers in the data strongly impact the non-simulated CRPS and MIS scores, our best model’s score falls within the value range of the simulation for all three metrics. As the windows with the lowest and highest share of non-zero pgms, we exemplarily discuss the simulation results for 2018 and 2023. Without any noise, CRPS values range from 0.007 to 0.172 and 0.010 to 0.252, respectively, dropping roughly linearly from “perfect” predictions to 0.1 accuracy. This means an average CRPS increase of 0.018 (2018) / 0.027 (2023) per 10% drop in accuracy. The more noise is added to the data, the smaller the change in CRPS becomes when varying the accuracy, with the changes less than half for the maximum level of noise in our simulation. Since accuracy in contrast to noise has an absolute floor, noise almost completely loses its impact on the score for very low levels of $\alpha$. Given that our best models are on average 0.007 points better than the best benchmark in the performance evaluation, this makes us confident that this means a real improvement in performance over the benchmarks rather than random variation. We observe very similar patterns for the MIS (Figure A1) and also for the IGN (Figure A2), with the caveat of the latter being much more sensitive to noise than CRPS and MIS and the noise curve for "perfect" accuracy appearing logarithmic rather than linear (see equation 2).
@@
Figure 4: Comparison of model-wide scores with scores based on grid cells grouped by individual countries for the three distributional metrics: a) CRPS, b) MIS, and c) IGN.
@@
Finally, having established that the small differences in the evaluation metrics are indeed meaningful, we want to understand where these differences come from. To do so, we match grid cells to countries based on a majority rule and evaluate our global model for each country individually. Country borders for assignment are taken from geoBoundaries (Runfola et al., 2020). Since this approach means strongly varying levels of zero-inflatedness across the countries, the corresponding scores also vary strongly and cannot be directly compared (Figure 4). Instead, we rank the performance of all models and benchmarks for each country and calculate the average ranks across all scenarios (Table 3). Based on the mean rank across all countries and years, our model only places fifth on the CRPS, fourth on the MIS, and third on the IGN. This can be explained by the fact that roughly half of all countries never experience any violence, for which our model gets heavily penalized in a rank-based evaluation due to its base assumption of a small but non-zero chance of violence anywhere. [^18] For both the CRPS and the IGN, any non-zero values in our samples mean that predictions will be scored marginally worse. Since the benchmarks generally yield all-zero samples in countries without prior violence, this means our model will be ranked 6th in such cases, while the benchmarks all receive a joint 1st place. To illustrate: Out of 80 countries covered by our model, 32 never see any fatalities across all yearly test windows, with a minimum of 39 no-violence countries in any individual window. However, when we restrict the analysis to country-year instances with any observed violence – arguably the cases in which predictive performance is most relevant – our model achieves the highest average rank compared to all benchmarks across all three distributional metrics. This likely indicates that our models’ predictions are relatively better for pgms where violence actually occurred.
@@ 5 Discussion
We demonstrate that it is possible to produce high-resolution conflict predictions with uncertainty estimates using a modeling architecture built from relatively simple components, which are able to beat a range of heuristic conflict benchmarks and competing approaches in a public prediction challenge. We further find that these performance gains arise from cases where violence occurs, and thus in situations in which prediction tools are likely most valuable to practitioners. With our evaluation, we highlight the need to not simply pursue marginal improvements in evaluation metrics, but to understand how they behave in the context of the given problem. While this is a useful first step towards actionable uncertainty estimates, the prediction intervals we produce are still fairly wide, with further research needed both to sharpen predictions and to reduce the sources of uncertainty.
@@
Building on our approach, data-related avenues for improvement naturally include the use of additional data sources, potentially in combination with multi-level modeling to combine data at different resolutions, while also accounting for the distinction between macro- and micro processes of conflict (Balcells and Justino, 2014; Fritz et al., 2024). As our results show, it is possible to combine multiple regional models without significantly losing performance. This opens up avenues to integrate data sources previously discarded due to coverage issues, for example collected by regional organizations such as ECOWAS. Moreover, data cleaning techniques such as outlier correction, applied both to the target and to the features, could reduce data-driven uncertainties. In addition to general improvements through better-suited methods, approaches such as selective classification (El-Yaniv and Wiener, 2010; Hüllermeier and Waegeman, 2021) and out-of-distribution detection (Yang et al., 2024; Gruber et al., 2025) could be used on the modeling side to reject predictions where the estimated uncertainty is too high for the predictions to be useful in practice.
@@
Finally, as little is known about the composition of uncertainty in conflict modeling, more research is needed to understand to what extent uncertainty is caused by data issues and to what extent violence is the consequence of inherently stochastic processes. Much of this question relates back to conflict datasets and the challenges of recording events globally as discussed above. For example, fine-grained local datasets may report many instances of violence not contained in the larger data collection efforts needed for forecasting systems (Bazzi et al., 2022). In combination with differences in inclusion criteria and coding procedures (Raleigh et al., 2023; Öberg and Yilmaz, 2025), this leads to a large number of non-overlapping events among conflict event datasets despite very similar sources (Donnay et al., 2019), and makes it hard to establish a ground truth for further examination. Addressing this issue would go a long way towards understanding and quantifying the individual components of uncertainty in conflict forecasting, and thus enable a more targeted approach to its reduction.